# Key exchange protocol based on circulant matrix action over congruence-simple semiring

### v1

@@ 3 Key exchange protocol and its security
First, we will see the computational cost of the key exchange Coste del set up
@@
La matriz $M\in Mat_{m}(R)$, tomamos como cota para los polynomios $h\in\mathbb{N}$, y $n$ elementos de la base. El coste del set up para el protocolo es a lo sumo $O(nH)$ operaciones matriciales, o $O(nhm^{3})$ operaciones en $R$
@@
Para cada elemento de la base, tenemos que calcular a lo sumo $h$ potencias, y dado que cada una se puede realizar con una multiplicacion de matrices de orden $m$, tenemos un coste de $O(2log_{2}(h)m^{3})$, tras eso, debemos realizar la suma, teniendo a lo sumo $H$ matrices que sumar, obteniendo $O(2log_{2}(h)m^{3}+2log_{2}(h))=O(2log_{2}(h)m^{3})$ operaciones en $R$. Puesto a que hay $n$ elementos en la base, el coste total es de $O(n2log_{2}(h)m^{3})$
Coste del algoritmo post set up
@@
Si $A\in Circ_{n}([0,k]$, entonces el coste de calcular $A(M_{i})_{i=1}^{n}$ es $O(n(2log_{2}k+n)m^{3})$ operaciones en $R$
@@
Cada producto o potencia tiene coste $m^{3}$. Tenemos que calcular potencias hasta $k$, lo cual tiene un coste de $O(2log_{2}k)$ potencias, dando lugar a $O(2log_{2}km^{3}$ en R, tras lo cual hay que multiplicar todas las matrices, por lo que añadimos $O(nm^{3})$ operaciones en $R$. Hay que repetir esto $n$ veces, obteniendo $O(n(2log_{2}k+n)m^{3})$ operaciones en $R$

### v2

@@ 3 Key Exchange Protocol and Its Security
We begin by analyzing the computational cost of the key exchange.
Setup Cost
@@
Let $M\in\text{Mat}_{m}(R)$ be a matrix over the ring $R$, with polynomials bounded by degree $h\in\mathbb{N}$, and let there be $n$ basis elements. Then, the setup cost of the protocol is at most $O(nH)$ matrix operations, or equivalently $O(nhm^{3})$ operations in $R$.
@@
For each basis element, we must compute up to $h$ powers of $M$. Each power can be computed using matrix multiplications of size $m\times m$, with a cost of $O(2\log_{2}(h)m^{3})$ in $R$. After computing the powers, we need to sum at most $H$ matrices, which contributes at most an additional $O(Hm^{2})$ operations; however, since matrix addition is negligible compared to multiplication, the overall cost per basis element remains $O(2\log_{2}(h)m^{3})$. Repeating this for all $n$ basis elements, the total setup cost becomes $O(n\log_{2}(h)m^{3})$ operations in $R$. ∎
Post-Setup Computational Cost
@@
Let $A\in\text{Circ}_{n}([0,k])$ be a circulant matrix with entries in the interval $[0,k]$. Then, the cost of computing $A(M_{i})_{i=1}^{n}$ is $O(n(2\log_{2}k+n)m^{3})$ operations in $R$.
@@
Each matrix multiplication or exponentiation involves $m^{3}$ operations. To compute powers up to $k$, we require $O(2\log_{2}k)$ multiplications, resulting in $O(2\log_{2}k\cdot m^{3})$ operations in $R$. After that, we must multiply the resulting matrices, contributing an additional $O(nm^{3})$ operations. Since this procedure is repeated for each of the $n$ components, the total cost is $O(n(2\log_{2}k+n)m^{3})$ operations in $R$. ∎